The ordered pairs in each exercise are for the same direct variation. Find each missing value. (1,y) and (2/3,-9)
Your help is greatly appreciated! THANK YOU!

Answer by stanbon(75887)

The ordered pairs in each exercise are for the same direct variation. Find each missing value. (1,y) and (2/3,-9)
---------------
y/1 = -9/(2/3)
---
y = -27/2
---
y = -13.5
-------
Cheers,
Stan H.
---------